OpenStudy (anonymous):

How do I solve 5^(-2x)=3?

OpenStudy (anonymous):

change of base says \[b^x-A\iff x=\frac{\ln(A)}{\ln (b)}\]

OpenStudy (anonymous):

make that \[b^x=A\iff x=\frac{\ln(A)}{\ln (b)}\]

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

you probably do not have log base 5 on your calculator, which is why you need the change of base formula. in one step you should write \[-2x=\frac{\ln(3)}{\ln(5)}\] and so \[x=\frac{\ln(3)}{-2\ln(5)}\]

OpenStudy (anonymous):

Awesome! Thanks--That's just what I needed :)

OpenStudy (anonymous):

yw